OpenStudy (anonymous):

Write the point-slope equation of the line through (-2, 1) with a slope of 3. y - 1 = 3(x + 2) y - 2 = 3(x + 1) y + 2 = 3(x - 1)

OpenStudy (anonymous):

y-y1 = m(x-x1) here y1 = 1 and x1 = -2 and m = 3 put these values in the above equation and get your answer

OpenStudy (anonymous):

okay